All times UTC.

Flats

12:14 Started continuous images in manual mode with Br_gam and 1.98s integration. Level ~5000.

12:17 Set to Ks and 1.98s. Level ~30,000. Switching back to Br_gram to flush detector. Some cirrus clouds on the horizon obscuring the Sun.

12:28 Counts in Br_gam down to 2000. Switch to K (won't be using Ks) and counts are 25,000.

12:33 Counts in Br_gam down to 1400. Switch to K and counts are 16,000. Setting to automatic mode in order to write files. Beginning 10 exposures at 1.98s integration (images #.0003 - #.0012).

12:39 Switching to H in manual mode, counts are 25,000. Using FeII filter to check sky level.

12:45 Back to H. Counts now 9,000 at 2s, using 4s int. time counts are 12,000 and falling fast, doing 10 exposures (#.0013 - #.0022). note -- for all previous images, image type was set to CALIB , now changing to FLAT

12:50 Switch to J, at 2s only 4000 counts. Now using 7s and taking 10 exposures (#.0023 - #.0032).

Pointing & Collimation

1:08 Attempt to do pointing correction, but collimation is well off (#.0033).

1:15 ptcorr is failing in an hselect command, looking for luci.0033 instead of luci.20100105.0033.fits. Fuyan solved it: hselect needed to be unlearned, because it had some parameter set that was causing problems.

1:24 now pointing model is updated. Way out of focus; T=2.6C. Applied a large Z offset and now collimating.

1:31 collimation finished. Seeing 0.5''.

J2329-0301 (Longslit)

1:33 Ran acquisition script for telluric standard. Waiting a few minutes for collimation. It seems to be well off.

1:43 Turned collimation off. Problem was that we were too close too limit for primary mirror. Applying an offset to it.

1:46 Rerunning the telluric acquistion script, now have a slit image since Wait for Collimation was off. Turning it back on and collimating. Still having problems. Some corrections removed and rerunning script. Still problems. Dave T. uploaded a different collimation model, then we applied a new pointing correction on a nearby bright, and now collimating on a Persson star.

2:46 Giving up after failing to collimate. Target was down to 20 deg elevation so sending the telescope to 70 deg elevation to attempt collimation again. All files #.0034 - .#0056 are junk related to testing pointing/collimation.

J0203+0012 (Longslit)

3:15 Acquistion script for faint target intregrated for 1 minute. But this is the time used for the slit image, so it should be short integration and then changed to long integration manually once the slit image is obtained.

A number of problems ensued. We applied a manual offset in the acquisition images in order to generate a sky-subtracted image of the field (the target is very faint). We then attempted to move the target into the slit and lost guiding. The guide star was near the edge of the patrol region, and our offsets moved it out of the field. We attempted to reacquire the target and move in a better direction but had continued difficulties with the telescope/guider, being looked into by LBT staff. We finally offset a total of 30'' and restarted the process of putting the target in the slit. For some reason, some image files (#75 and #77) were not visible in skycat, so we had to take extra images in order to select the target.

4:45 After those several offsets to the north (-Y in DETXY) the target is at y=780. Applying an offset (dx,dy)=(62.00,7.02). The slit image is #73, the final acquisition images are #76 and #78.

4:51 Starting quasar observation script. The trace is barely visible in the first images, at y~850 in #79, and y~800 in #80. Some passing cirrus. Seeing 0.7''. This first set is #.0079 - #.0086 with DIT=150 NDIT=1 and 8 offsets.

5:17 Due to how faint the object is, we changed the script to DIT=150 NDIT=2 and ran the script again. We also changed the first offset to "0 -5" in order to have the target land where it began on image #79. This set is #87 - #94. Still passing cirrus.

5:22 first image #87 we temporarily lost guiding due to clouds. Definitely hazy.

5:40 gave up. last image is #88. background went up, guiding difficult. Closing dome for safety, cloud cover is light but across full sky. Waiting a while for it to clear.

J0900+2234 (Longslit)

8:12 Reopening dome.

8:15 Assuming pointing is still okay, going to Persson star for collimation. Well out of focus. Collimation was fairly quick. Sky mostly clear with some high, light cirrus.

8:24 Acquiring target. Seeing 0.5''. Slit image is #90. Script uses an offset from nearby star. Offset star found at (1016,1118), moved to (1026,618) using skycat which is an offset of (1.17',-60.06'').

8:57 Starting science observations. When the offset was applied to go to the science target, the PA in the Current Target part of Telescope Control went to 0. But under Pointing it says 89.67, which is what it should be according to the script. Currently unknown whether instrument is at correct PA, taking exposures to see if object is visible. Series begins at #97.

9:15 Stopped at #99 because unsure whether target is in slit. Moving mirror in to take a look at field. #100 and #101 are images of the field. Pointing and rotation look good so continuing with spectra.

9:28 restarted science script. Images #102-106.

10:00 moving to telluric. More passing cirrus. Telluric clearly in slit during acquisition image (#107) so taking observations.

10:05 telluric image #108.

10:08 lamp exposures #109-#113.

J0959+0206 (COSMOS4)

10:27 attempting acquisition. slit image #114. acquisition using #120 and #118, the star north of the target that should also be in the slit is at 1025,880. Moved it into slit.

10:46 beginning science observation at #121.

11:00 aborted because guide probe is clearly vignetting field (emission in spectrum). the selected guide star is too close to Lucifer FOV. Tried to move to another guide star and it failed because it would have moved the probe to an invalid position. Tried another guide star and failed again. Then noticed that the PA of the instrument had arbitrarily moved to 243 when it should have been at 184 (where it was when we began). We changed the guide star in the script to a better one and then reran the acquisition script to reset the position.

11:16 reacquiring with #126/#125. Only need a small offset (0.88,-9.01). No emission from probe and guiding is stable.

11:19 beginning science observation at #128.

12:10 interfaces are frozen on all local computers due to network problem. unclear whether lucifer is still taking data.

1:33 still trying to resolve network issues. seems to be problem with NFS server.

-- FuyanBian - 03 Jan 2010
Topic revision: r8 - 05 Jan 2010, IanMcGreer
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback